# Efficient Neural Video Representation via Structure-Preseving Patch Decoding

### v1

@@ Abstract
Implicit Neural Representations (INRs) have attracted significant interest for their ability to model complex signals by mapping spatial and temporal coordinates to signal values. In the context of neural video representation, several decoding strategies have been explored to balance compactness and reconstruction quality, including pixel-wise, frame-wise, and patch-wise methods. Patch-wise decoding aims to combine the flexibility of pixel-based models with the efficiency of frame-based approaches. However, conventional uniform patch division often leads to discontinuities at patch boundaries, as independently reconstructed regions may fail to form a coherent global structure. To address this limitation, we propose a neural video representation method based on Structure-Preserving Patches (SPPs). Our approach rearranges each frame into a set of spatially structured patch frames using a PixelUnshuffle-like operation. This rearrangement maintains the spatial coherence of the original frame while enabling patch-level decoding. The network learns to predict these rearranged patch frames, which supports a global-to-local fitting strategy and mitigates degradation caused by upsampling. Experiments on standard video datasets show that the proposed method improves reconstruction quality and compression performance compared to existing INR-based video representation methods.
@@ I Introduction
Implicit Neural Representations (INRs) have emerged as a compact, resolution-independent alternative to conventional explicit representations for modeling complex signals. By expressing data as continuous functions of spatial and temporal coordinates through neural networks, INRs significantly reduce memory usage. They also support desirable properties such as infinite-resolution rendering and smooth interpolation. These advantages make them well-suited for modeling fine-grained signals in domains such as novel view synthesis [1, 2, 3, 4], image representation [5, 6, 7, 8], and increasingly, video representation [9, 10, 11].
@@
Figure 1: Comparison of decoding strategies for neural video representation. (a) Coordinate-based: predicts pixels from spatio-temporal coordinates. (b) Frame-based: generates entire frames from embeddings. (c) Patch-based: reconstructs frames from uniformly divided patches. (d) SPP-based (ours): rearranges pixels into spatially aligned patches for structure-preserving, global-to-local reconstruction.
INR-based methods for neural video representation can be categorized by their decoding granularity: coordinate-based (pixel-wise) [12, 13], frame-based (image-wise) [9, 10, 11], and patch-based approaches [14, 15]. Each approach presents a different trade-off between flexibility, computational cost, and structural consistency. Coordinate-based approaches predict individual pixel values based on spatial and temporal inputs, offering unmatched resolution flexibility and the ability to capture intricate local variations. However, they are often computationally intensive and limited in capturing global context due to their inherently localized nature. Frame-based approaches, such as Neural Representations for Videos (NeRV) [9], mitigate this issue by predicting entire frames from compact embeddings, such as time indices or learned content vectors. These models benefit from faster inference and improved global structure modeling, but tend to struggle with high-frequency detail due to upsampling artifacts and spectral bias (favoring low-frequency components) [16, 17, 18, 19]. Patch-based methods lie between these extremes. By reconstructing each frame from smaller subregions, they aim to combine the detail-preserving nature of pixel-wise decoding with the structural coherence of frame-wise decoding. Nonetheless, conventional patch partitioning via uniform spatial division can disrupt continuity across patch boundaries. Independently reconstructed patches often fail to align perfectly, resulting in visible seams and reduced visual quality. This limitation motivates the exploration of patch-based methods that maintain global spatial coherence without sacrificing local detail fidelity.
@@
Figure 2: Toy experiment comparing the fitting performance of one-dimensional signals. Top row (left to right): ground-truth signal, point-wise fitting, and whole-wise fitting. Bottom row (left to right): training loss curves, patch-wise fitting, and the proposed Structure-Preserving Patch (SPP)-wise fitting. While conventional patch-wise approaches suffer from boundary discontinuities, the proposed SPP-wise approach maintains structural continuity and enables accurate local fitting.
To overcome these limitations, we propose the Structure-Preserving Patch (SPP) method for neural video representation. Instead of independently decoding spatially divided patches, our method rearranges each frame into multiple sub-images through a pixel redistribution operation inspired by PixelUnshuffle. This operation preserves the relative spatial structure of the original frame across all patches. The model is trained to predict these structurally aligned patch images, enabling it to first capture global structure and then refine local content. This global-to-local fitting process reduces the risk of discontinuities at patch boundaries and leads to improved visual quality. We validate the effectiveness of our method through experiments on several benchmark video datasets. The results show that our structure-aware patch decoding improves reconstruction quality and compression performance compared to prior INR-based video representation models.
@@ II-A Implicit Neural Representation
where $x$ is a continuous coordinate, $y$ is the signal value at that location, and $\theta$ denotes the learnable parameters of the network. The function $\Phi_{\theta}$ is trained to fit a given target signal through optimization. A seminal work in this field is Neural Radiance Fields (NeRF) [1], which models a 3D scene as a radiance field defined over spatial coordinates and viewing directions. NeRF demonstrated that photo-realistic view synthesis is possible by optimizing a fully connected network to predict color and density values along camera rays. Although computationally demanding, NeRF highlights the potential of neural functions to model complex continuous structures without relying on explicit mesh or voxel representations. To enhance the capability of INRs in modeling high-frequency details, SIREN [12] introduced a sinusoidal-activated network capable of modeling high-frequency content with improved gradient flow. By leveraging periodic activation functions, SIREN reduces the spectral bias inherent in ReLU-based networks and improves the network’s ability to capture fine details and textures. These foundational works have inspired the application of INRs to spatio-temporal domains, including video data. Their success shows that compact and continuous neural representations can be extended from static to dynamic signals, motivating recent research into INR-based video modeling.
Figure 3: Overview of the proposed Structure-Preserving Patchs (SPPs) based video representation framework. (a) Overall architecture. An input frame $I_{t}$ is encoded into a frame-level embedding $z_{t}$, which is then decoded into patch images $I_{t,i}$ using the temporal index $t$ and patch index $i$. The full frame is reconstructed by spatially rearranging the decoded patches. (b) Decoder details. Temporal and patch indices are embedded using positional encodings and MLPs. Early decoder layers are conditioned on $t$ to capture global structure, while later layers use $i$ to refine local patch details. (c) Patch construction via SPPs. Each frame is rearranged into multiple patch images (e.g., $I_{t,0}$, $I_{t,1}$, $I_{t,2}$, $I_{t,3}$) that preserve spatial consistency. This design supports a smooth global-to-local reconstruction process.
@@ II-B Neural Video Representation
In addition, patch-wise and hybrid methods have been proposed to improve reconstruction quality and scalability. For instance, PS-NeRV [14] reconstructs videos on a patch-by-patch basis instead of generating full frames in one shot. This patch-wise process offers flexibility in handling higher resolutions and can distribute computational load, albeit with the risk of patch boundary artifacts. To better capture video dynamics, DNeRV [22] adopts a dual-stream architecture. One stream models the static background, while the other captures inter-frame differences, allowing motion to be represented implicitly. Several other approaches introduce hierarchical model structures [23, 24, 25], incorporating explicit residual connections [26, 27], or emphasizing the preservation of frequency-domain characteristics [28, 29, 30]. Each of these innovations addresses specific shortcomings of basic INR video models, such as temporal flicker and poor high-frequency reconstruction. Our work aligns with this line of research, focusing on patch-based representation. We specifically aim to preserve global image structure in a patch-wise model to avoid the common pitfalls of patch division.
@@ III-A Overview
We propose a neural video representation method based on Structure-Preserving Patches (SPPs), as illustrated in Fig. 3. In this framework, each frame is decoded at the patch level. Unlike conventional patch-based methods, however, the patches are constructed to retain the original spatial layout of the frame. The key idea is to avoid treating patches as independent tiles and to generate patch images in a way that preserves the layout of the frame. These patch images tend to share similar structures and contain redundant information, which the network can exploit to improve representation efficiency. The decoding process is designed to follow a global-to-local fitting strategy. The model first learns a coarse global representation of the frame and then progressively refines local details within each patch. This hierarchical fitting process improves the network’s ability to reconstruct both global context and fine details. By preserving spatial continuity across patch boundaries, our method mitigates the quality degradation commonly observed in naive patch-based upsampling or reconstruction.
@@ III-B Motivation
To highlight the importance of the global-to-local strategy, we conduct a simple one-dimensional signal-fitting experiment (Fig. 2). We compare four strategies: (1) Point-wise, where the signal is fitted at each coordinate independently; (2) Patch-wise, where the signal is divided into fixed-length segments and each segment is fitted separately; (3) Whole-wise, where the entire signal is fitted as a single input; and (4) SPP-wise, our proposed approach, which fits a rearranged version of the signal based on fixed-position samples, mimicking the structure-preserving patch mechanism. The results show that the whole-wise and SPP-wise strategies achieve superior fitting accuracy by preserving the overall structure of the signal. In contrast, point-wise and patch-wise methods prioritize local fitting and fail to maintain global consistency. These observations suggest that modeling the global structure first, followed by local refinement, is essential for accurate signal reconstruction. This insight forms the foundation of our global-to-local decoding framework used for video representation.
@@
Figure 4: Visualization of reconstructed frames from different sequences. From top to bottom: “drift-straight” and “mallard-fly” from the DAVIS dataset, and sequences #10 and #12 from the MCL-JCV dataset. Our structure-preserving patch (SPP) based method maintains sharp edges and consistent spatial structure across frames, demonstrating effective reconstruction quality across diverse scenes.
III-C Structure-Preseving Patch Decoding
@@
Our framework predicts each video frame at the patch level using a neural network that follows a global-to-local decoding strategy, as illustrated in Fig.3. For each input frame $I_{t}$ at time $t$, we divide it conceptually into $N$ spatially structured patches. To implement this, we apply a pixel rearrangement operation, similar to PixelUnshuffle, which transforms the frame into $N$ patch images: $I_{t,0},I_{t,1},\dots,I_{t,N-1}$. This rearrangement preserves the original spatial relationships across patches. The encoder, implemented using ConvNeXt blocks [31, 11], first processes the full frame $I_{t}$ to produce a compact frame-level embedding $z_{t}$. This embedding is passed to a decoder, which generates each patch image $I_{t,i}$ under the influence of two indices: the temporal index $t$ and the patch index $i$. In the decoder, early layers are modulated only by $t$, allowing the model to capture the global structure shared across all patches in the frame. Later layers are additionally conditioned on $i$, enabling the network to refine local details specific to each patch. Once all patch outputs $I_{t,i}$ are obtained, we reconstruct the full-resolution frame $\hat{I}_{t}$ by spatially rearranging them. This hierarchical conditioning mechanism allows the decoder to first build a coherent global structure, then specialize to local content, resulting in improved reconstruction quality. In addition, the structure-preserving nature of our patch generation helps avoid boundary discontinuities typically associated with naive patch-based decoding. The decoder is composed of multiple blocks $B_{n}$, each of which integrates the SNeRV block and the TAT Residual Block introduced in BoostingNeRV [21].
@@ III-D Loss Function
[…]
@@ IV Experiment
We evaluate the effectiveness of the proposed SPP-based video representation method across four benchmark datasets: Bunny [32], DAVIS [33], MCL-JCV [34], and UVG [35]. Our method is compared with existing implicit neural video representation baselines, including NeRV, HNeRV, and Boosting-HNeRV (“Boost”). The Bunny dataset consists of a single sequence with a resolution of $720\times 1280$ and 132 frames. To ensure stable training, the frames are cropped to $640\times 1280$. This adjustment is made because HNeRV failed to converge when trained on high-resolution input, as observed in the “Bosphorus” sequence in Table III. The DAVIS dataset consists of 50 natural video sequences with resolution $1080\times 1920$ (cropped $640\times 1280$). Each sequence is relatively short (25 to 104 frames). Similarly, MCL-JCV contains 30 sequences at $1080\times 1920$ (cropped to $640\times 1280$), with lengths ranging from 120 to 150 frames. The UVG dataset comprises 7 long sequences at full HD resolution ($1080\times 1920$), each containing either 300 or 600 frames.
@@
We adjust the stride configuration in the decoder to match the spatial resolution: the stride list is set to [5, 4, 2, 2, 2] for $640\times 1280$, and [5, 3, 2, 2] for $1080\times 1920$. The number of patches is fixed at $N=4$, and we empirically set the loss weighting coefficients to $\alpha=42$ and $\beta=18$. The default model size is 1.5M parameters, except for the UVG dataset at $1080\times 1920$ resolution, where a 3.0M parameter model is used. For quantitative evaluation, we use Peak Signal-to-Noise Ratio (PSNR), Multi-Scale Structural Similarity (MS-SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). In the video compression setting, we use the UVG dataset with a resolution of $1080\times 1920$ and subsample each sequence to 60 or 120 frames. Each video is encoded into multiple models of varying sizes by adjusting the network width. We apply quantization and entropy coding to these models to simulate real-world compression, enabling rate-distortion (RD) analysis.
@@
Quantitative results on DAVIS and MCL-JCV are summarized in Tables I and II, respectively. These tables show the average reconstruction quality across each datasets. Table III reports PSNR and MS-SSIM scores for each sequence in the UVG dataset. Our method consistently achieves higher scores in most sequences, demonstrating superior reconstruction quality compared to baseline methods. In particular, Table IV presents PSNR values at each training epoch for the Bunny dataset, demonstrating not only improved reconstruction performance, but also faster convergence compared to competing methods. Qualitative comparisons are shown in Fig. 4, which visualizes representative frames from DAVIS and MCL-JCV. Our method better preserves sharp edges and spatial coherence across a variety of scenes. Finally, Fig. 5 presents the rate-distortion curves obtained from UVG. The proposed method achieves superior compression performance relative to the baselines. Note that HNeRV is excluded from the average curve due to repeated convergence failures during training.

### v2

# Structure-Preserving Patch Decoding for Efficient Neural Video Representation

@@ Abstract
Implicit neural representations (INRs) are the subject of extensive research, particularly in their application to modeling complex signals by mapping spatial and temporal coordinates to corresponding values. When handling videos, mapping compact inputs to entire frames or spatially partitioned patch images is an effective approach. This strategy better preserves spatial relationships, reduces computational overhead, and improves reconstruction quality compared to coordinate-based mapping. However, predicting entire frames often limits the reconstruction of high-frequency visual details. Additionally, conventional patch-based approaches based on uniform spatial partitioning tend to introduce boundary discontinuities that degrade spatial coherence. We propose a neural video representation method based on Structure-Preserving Patches (SPPs) to address such limitations. Our method separates each video frame into patch images of spatially aligned frames through a deterministic pixel-based splitting similar to PixelUnshuffle. This operation preserves the global spatial structure while allowing patch-level decoding. We train the decoder to reconstruct these structured patches, enabling a global-to-local decoding strategy that captures the global layout first and refines local details. This effectively reduces boundary artifacts and mitigates distortions from naive upsampling. Experiments on standard video datasets demonstrate that our method achieves higher reconstruction quality and better compression performance than existing INR-based baselines.
@@ I Introduction
Implicit Neural Representations (INRs) have gained attention as compact and resolution-independent alternatives to conventional explicit representations for modeling complex signals. By parameterizing data as continuous functions of spatial and temporal coordinates using neural networks, INRs significantly reduce memory usage while enabling smooth interpolation and reconstruction of signals at arbitrary resolutions. These properties have made INRs effective in areas ranging from novel-view synthesis [1, 2, 3, 4] to static image representation [5, 6, 7, 8] and, more recently, to video representation [9, 10, 11].
@@
Figure 1: Comparison of decoding strategies for neural video representation. (a) Coordinate-based: predicts pixel color from spatial and temporal coordinates. (b) Frame-based: generates entire frames from embeddings. (c) Patch-based: reconstructs frames from uniformly divided patches. (d) SPP-based (ours): arranges pixels into spatially aligned patches for structure-preserving, global-to-local reconstruction.
Existing INR-based video representation methods can be categorized into three primary types, depending on their decoding granularity: (1) coordinate-based approaches [12, 13], which predict each pixel from its spatiotemporal coordinates; (2) frame-based approaches [9, 10, 11], which generate entire frames from compact input; and (3) patch-based approaches [14, 15], which reconstruct frames by predicting spatially partitioned patches. While coordinate-based approaches provide fine-grained spatial control and support resolution scalability, their independent pixel-wise processing limits the ability to capture global scene structure and incurs high computational costs. To address these issues, frame-based methods generate entire frames from compact temporal or content embeddings, enabling faster inference and more effective modeling of global structure. However, they often struggle to preserve high-frequency components due to upsampling artifacts and the spectral bias of neural networks, which favors low-frequency components [16, 17, 18, 19]. In addition, patch-based decoding lies between these extremes. Reconstructing each frame from smaller spatial regions combines the local detail fidelity of pixel-based methods with the global structural modeling of frame-based methods. However, uniform spatial partitioning often leads to discontinuities across patch boundaries, as independently reconstructed regions lack alignment. As a result, visual artifacts such as seams and diminished perceptual quality occur.
@@
Figure 2: Toy experiment comparing the fitting performance of one-dimensional signals. Top row (left to right): ground-truth signal, point-wise fitting, and global fitting. Bottom row (left to right): training loss curves, segment-wise fitting, and the proposed Structure-Preserving Patch (SPP)-style fitting. While conventional segment-wise approaches suffer from boundary discontinuities, the proposed SPP-style approach maintains structural continuity and enables accurate local fitting.
In order to overcome these limitations, we propose a Structure-Preserving Patches (SPPs) decoding method for neural video representation. Rather than dividing frames into spatially disconnected patches, our method applies a pixel arrangement operation, similar to PixelUnshuffle, to generate patch images that preserve the original spatial layout. This pixel-based splitting maintains the relative spatial relationships among patches, enabling the network to first capture global scene structure and later focus on refining local details. To implement this, the decoder is structured so that early layers are guided by the content embedding and temporal index to model global structure, with later layers incorporating the patch index to specialize in local refinement. This global-to-local reconstruction strategy reduces patch boundary artifacts and enhances visual coherence. We validate our approach on multiple benchmark video datasets. The experimental results demonstrate that our method achieves superior reconstruction quality and compression performance compared to existing INR-based video representation methods.
@@ II-A Implicit Neural Representation
where $x$ is a continuous coordinate, $y$ is the signal value at that location, and $\theta$ denotes the learnable parameters of the network. The network $\Phi_{\theta}$ is optimized to fit a given target signal. A seminal work in this field is Neural Radiance Fields (NeRF) [1], which models a 3D scene as a radiance field defined over spatial coordinates and viewing directions. NeRF demonstrated that photo-realistic view synthesis is possible by optimizing a fully connected network to predict color and density values along camera rays. Although computationally demanding, NeRF highlights the potential of neural functions to model complex continuous structures without relying on explicit mesh or voxel representations. By leveraging periodic activation functions, SIREN [12] reduces the spectral bias inherent in ReLU-based networks and improves the network’s ability to capture fine details and textures. These works show that compact and continuous neural representations have inspired the application of INRs to spatio-temporal domains, including video data.
Figure 3: Overview of the proposed Structure-Preserving Patches (SPPs) based video representation framework. (a) Overall architecture. An input frame $I_{t}$ is encoded into a frame-level embedding $z_{t}$, which is then decoded into patch images $I_{t,i}$ using the temporal index $t$ and patch index $i$. The full frame is reconstructed by spatially rearranging the decoded patches. (b) Decoder details. Temporal and patch indices are embedded using positional encodings and MLPs. Early decoder layers are conditioned on $t$ to capture global structure, while later layers use $i$ to refine local patch details. (c) Patch construction via SPPs. Each frame is split into multiple patch images (e.g., $I_{t,0}$, $I_{t,1}$, $I_{t,2}$, $I_{t,3}$) that preserve spatial consistency. This design supports a smooth global-to-local reconstruction process.
@@ II-B Neural Video Representation
Recent works have also explored alternative decoding strategies to improve reconstruction quality and scalability. For example, PS-NeRV [14] reconstructs video frames on a patch-wise basis, rather than decoding full frames at once. This design provides flexibility in handling high-resolution content and facilitates parallel computation, but often introduces visible artifacts along patch boundaries due to the lack of spatial continuity. To better capture video dynamics, DNeRV [22] adopts a dual-stream architecture. One stream models the static background, while the other captures inter-frame differences, allowing motion to be represented implicitly. Several other approaches introduce hierarchical model structures [23, 24, 25], incorporating explicit residual connections [26, 27], or emphasizing the preservation of frequency-domain characteristics [28, 29, 30], in order to mitigate issues such as temporal flickering or poor reconstruction of fine details. Our approach follows this research direction, focusing on patch-based representation. Distinct from prior patch-based methods, we explicitly aim to preserve global spatial structure within the decoding process to mitigate boundary artifacts and improve perceptual consistency.
@@ III-A Overview
We propose a neural video representation method based on Structure-Preserving Patches (SPPs), as illustrated in Fig. 3. In this method, each video frame is decoded from a set of spatially aligned patch images. To retain the spatial layout of the original frame, we employ a deterministic pixel-based splitting operation, similar to PixelUnshuffle, instead of dividing frames into independently processed tiles as in conventional patch-based methods. This operation allows the network to exploit redundancy across patch images that share a coherent spatial arrangement, enhancing its ability to model global scene structure. To further improve reconstruction quality, we design the decoder to follow a global-to-local fitting strategy. In the early layers, the decoder models global structure using a content embedding and temporal index. Subsequently, patch indices are incorporated into the later layers to refine local details within each patch. This hierarchical decoding process enables accurate reconstruction of both global layout and high-frequency visual details.
@@ III-B Motivation
To highlight the importance of the global-to-local strategy, we conduct a simple one-dimensional signal-fitting experiment (Fig. 2). We compare four strategies: (1) Point-wise fitting, where the signal is fitted at each coordinate independently; (2) Segment-wise fitting, where the signal is divided into fixed-length segments, and each segment is fitted independently; (3) Global fitting, where the entire signal is fitted as a single input; and (4) SPP-style fitting, our proposed approach, which fits an arranged version of the signal based on fixed position samples, mimicking the structure-preserving patch mechanism. The results show that the global and SPP-style strategies achieve superior fitting accuracy by preserving the overall structure of the signal. In contrast, point-wise and segment-wise methods prioritize local fitting and fail to maintain global consistency. These observations suggest that modeling the global structure first, followed by local refinement, is essential for accurate signal reconstruction. This insight forms the foundation of our global-to-local decoding framework used for video representation.
@@
Figure 4: Visualization of reconstructed frames from different sequences. From top to bottom: “drift-straight” and “mallard-fly” from the DAVIS dataset, and sequences #10 and #12 from the MCL-JCV dataset. Our Structure-Preserving Patches (SPPs) based method maintains sharp edges and consistent spatial structure across frames, demonstrating effective reconstruction quality across diverse scenes.
III-C Structure-Preserving Patch Decoding
@@
Our framework predicts each video frame at the patch level using a neural network that follows a global-to-local decoding strategy, as illustrated in Fig.3. For each input frame $I_{t}$ at time $t$, we divide it conceptually into $P$ spatially structured patches. To implement this, we apply a pixel arrangement operation, similar to PixelUnshuffle, which transforms the frame into $P$ patch images: $I_{t,0},I_{t,1},\dots,I_{t,P-1}$. The encoder, implemented using ConvNeXt blocks [31, 11], first processes the full frame $I_{t}$ to produce a compact frame-level embedding $z_{t}$. This embedding is passed to a decoder, which generates each patch image $I_{t,i}$ conditioned on two indices: the temporal index $t$ and the patch index $i$. In the decoder, early layers are modulated only by $t$, allowing the model to capture the global structure shared across all patches in the frame. Later layers are additionally conditioned on $i$, enabling the network to refine local details specific to each patch. Once all patch outputs $I_{t,i}$ are obtained, we reconstruct the full-resolution frame $\hat{I}_{t}$ by spatially rearranging them. This hierarchical conditioning mechanism allows the decoder to first build a coherent global structure, then specialize to local content, resulting in improved reconstruction quality. In addition, the structure-preserving nature of our patch generation helps avoid boundary discontinuities typically associated with naive patch-based decoding. The decoder is composed of multiple blocks $B_{n}$, each of which integrates the SNeRV block and the TAT Residual Block introduced in BoostingNeRV [21].
@@ III-D Loss Function
where $x_{i}$ and $\hat{x}_{i}$ denote the ground-truth and predicted patch images, respectively. The term $\mathcal{L}_{1}$ is the standard L1 pixel loss, and $\mathcal{L}_{\text{MS-SSIM}}$ is a multi-scale structural similarity loss, which promotes perceptual quality. The coefficients $\alpha$ and $\beta$ balance their relative contributions. In addition, The frequency domain term $\mathcal{L}_{\text{freq}}$ is computed by applying the Fast Fourier Transform (FFT) to both images and measuring their L1 difference.
In our SPP-based decoding scheme, patch-specific variations are adjusted only at the later layers of the decoder. This localized modulation makes it difficult for the model to learn patches that differ significantly from their neighbors. To address this, we introduce an adaptive patch weighting factor $w_{i}$ as follows:
@@ IV Experiment
We evaluate the effectiveness of the proposed SPP-based video representation method across four benchmark datasets: Bunny [32], DAVIS [33], MCL-JCV [34], and UVG [35]. The Bunny dataset consists of a single sequence with a resolution of $720\times 1280$ and 132 frames. To ensure stable training, the frames are cropped to $640\times 1280$. This adjustment is made because HNeRV failed to converge when trained on high-resolution input, as observed in the “Bosphorus” sequence in Table III. The DAVIS dataset consists of 50 natural video sequences with resolution $1080\times 1920$ (cropped $640\times 1280$). Each sequence is relatively short (25 to 104 frames). Similarly, MCL-JCV contains 30 sequences at $1080\times 1920$ (cropped to $640\times 1280$), with lengths ranging from 120 to 150 frames. The UVG dataset comprises 7 long sequences at full HD resolution ($1080\times 1920$), each containing 300 or 600 frames. Our method is compared with existing implicit neural video representation baselines, including NeRV, HNeRV, and Boosting-HNeRV (“Boost”).
@@
We adjust the stride configuration in the decoder to match the spatial resolution: the stride list is set to [5, 4, 2, 2, 2] for $640\times 1280$, and [5, 3, 2, 2] for $1080\times 1920$. The number of patches is fixed at $P=4$, and we empirically set the loss weighting coefficients to $\alpha=42$ and $\beta=18$. The default model size is 1.5M parameters, except for the UVG dataset at $1080\times 1920$ resolution, where a 3.0M parameter model is used. For quantitative evaluation, we use PSNR, MS-SSIM, and LPIPS. In the video compression setting, we use the UVG dataset with a resolution of $1080\times 1920$ and subsample each sequence to 60 or 120 frames. Each video is encoded into multiple models of varying sizes by adjusting the network width. We apply quantization and entropy coding to these models, enabling rate-distortion (RD) analysis.
@@
Quantitative results on DAVIS and MCL-JCV are summarized in Tables I and II, respectively. These tables show the average reconstruction quality across each dataset. Table III reports PSNR and MS-SSIM scores for each sequence in the UVG dataset. Our method consistently achieves higher scores in most sequences, demonstrating superior reconstruction quality compared to baseline methods. In addition, Table IV presents PSNR values at each training epoch for the Bunny dataset, demonstrating not only improved reconstruction performance but also faster convergence compared to competing methods. Qualitative comparisons are shown in Fig. 4, which visualizes representative frames from DAVIS and MCL-JCV. Our method better preserves sharp edges and spatial coherence across a variety of scenes. Finally, Fig. 5 presents the rate-distortion curves obtained from UVG. Note that HNeRV is excluded from the average curve due to repeated convergence failures during training. The proposed method demonstrates superior compression performance relative to baseline INR methods. However, it does not achieve the same performance as traditional video compression technology, HEVC [36], and there is still potential for improvement in model compression.